OpenStudy (anonymous):

What are the solutions? 1/2x^2+ 2x + 3 =0 I know how to do trinomial factoring, I just can't seem to find the solutions though since the 1/2x^2 is kinda weird.

OpenStudy (anonymous):

Someone help pleasee!

OpenStudy (anonymous):

It has imaginary roots.

OpenStudy (anonymous):

Someone please, I don't get it

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x ^{2}+2x+3=0\] \[2\left( \frac{ 1 }{ 2 }x ^{2}+2x+3=0\ \right)\] \[x ^{2}+4x+6=0\]

OpenStudy (anonymous):

now use quadatic func

OpenStudy (anonymous):

What is that? how do I use quadratic function? Can you please explain?

OpenStudy (anonymous):

Are the solutions where the parabola touches the x-axis?

OpenStudy (anonymous):

The Quadratic Formula. look it up on youtube

OpenStudy (anonymous):

Ok, thanks, I am doing that now.

OpenStudy (anonymous):

So I got  Is that my solution, or does that simplify further?